The following is the information from the national income accounts for a hypothetical country:
GDP Rs. 6000.00
Gross Investment 800.00
Net Investment 200.00
Consumption 4000.00
Govt. purchases of goods & services 1100.00
Govt. Budget Surplus 30.00
What is
a) NDP
b) Net exports
c) Govt. taxes minus transfers
c) Disposable personal income
e) Personal Saving.
1
Expert's answer
2019-05-08T10:03:22-0400

a) NDP = Y - (Gross Investment - Net Investment) = 6000 - (800 - 200) = 5400.

b) Net exports NX = Y - G - I - C = 6000 - 1100 - 200 - 4000 = 700.

c) Govt. taxes minus transfers T = 1100 + 30 = 1130.

d) Disposable personal income is Yd = Y - T = 6000 - 1130 = 4870.

e) Personal Saving = Yd - C = 4870 - 4000 = 870.